OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the points (2,5) and (4,3)? A. y = x + 7 B. y = -x + 7 C. y = 2x + 5 D. y = -2x + 5

OpenStudy (anonymous):

would i do y2-y1/x2-x1 for this type of problem?

OpenStudy (shamil98):

Yes, you would find the slope first.

OpenStudy (shamil98):

3-5 / 4 - 2 = -1 Then plug in the values of either equation to get the equation of those two coordinates into point-slope form. y - y1 - m(x- x1)

OpenStudy (shamil98):

y-y1 = m(x-x1) *

OpenStudy (anonymous):

y=-x+7?

OpenStudy (shamil98):

y - 5 = -1(x-2) y - 5 = -x + 2 y = -x + 7 It says y1 and x1 so always use the first pair of coordinates , i'm rusty with algebra..

OpenStudy (shamil98):

Yeah.

OpenStudy (anonymous):

thanks!!